Given the following equations:


y = 8t2 – 6t – 11

and

x = 2t – 7


find y as a function of x and then give the value for f(8).

That is, solve for y in terms of x and replace y with the "f(x)" notation.

Found 2 solutions by rothauserc, Fombitz:
Answer by rothauserc(4718)

solve second equation for t
2t = x+7
t = (x+7) / 2
substitute for t in first equation
y = 8((x+7)/2)^2) -6((x+7) / 2) - 11
y = 8(x^2+14x+49)/4 -3(x+7) -11
y = 2x^2 +28x+98 -3x -21 -11
y = 2x^2 +25x +66
f(8) = 2(8^2) +25*8 +66
f(8) = 394

Answer by Fombitz(32388)